# Long-lived quasinormal modes, shadows and particle motion in four-dimensional quasi-topological gravity

### 11 Mar 2026 v1

@@ Abstract
We investigate massive scalar perturbations and several characteristics of particle motion in the spacetime of regular black-holes arising in four-dimensional quasi-topological gravity. Quasinormal modes are computed using high-order WKB approximations with Padé resummation and verified through time-domain integration. For moderate values of the scalar-field mass the time-domain profiles confirm the WKB results with excellent accuracy. As the mass increases, the damping rate decreases substantially, indicating the approach to the quasi-resonant regime of long-lived modes. For sufficiently large masses the late-time signal becomes dominated by oscillatory power-law tails, which mask the quasi-resonant mode in the time-domain profile. In addition, we analyze photon motion and circular geodesics, including the photon-sphere radius, shadow size, Lyapunov exponent, and ISCO characteristics. These quantities exhibit only moderate deviations from their Schwarzschild values, unlike the Hawking temperature of the black-hole.
@@ I Introduction
One of the recent findings in this area is the construction of regular black holes in four-dimensional non-polynomial quasi-topological gravity Borissova and Carballo-Rubio (2026); Bueno et al. (2026). Within this framework, black-hole solutions can be obtained analytically. The resulting geometries are asymptotically flat and free of curvature singularities, reproducing well-known regular metrics without introducing additional matter sources. A regularization parameter usually controls the deviation from the Schwarzschild spacetime and is typically localized in the near-horizon region.
Understanding how such geometrically regular black holes manifest themselves through observable phenomena requires analyzing both wave dynamics and particle motion in these spacetimes. One of the most important probes of black-hole geometry is provided by quasinormal modes (QNMs), which describe the characteristic oscillations of perturbed black-holes Konoplya and Zhidenko (2011); Bolokhov and Skvortsova (2025d); Kokkotas and Schmidt (1999). These modes dominate the ringdown stage of gravitational-wave signals produced in black-hole mergers and therefore play a key role in extracting information about the underlying black-hole spacetime geometry Abbott and others (2016, 2017, 2020). At the same time, properties of geodesic motion, such as the angular velocity at the innermost stable circular orbit (ISCO), the Lyapunov exponent governing the instability of photon orbits, the binding energy of particles, and the radius of the black-hole shadow, provide complementary observational diagnostics that directly probe the structure of the gravitational field in the strong-curvature regime.
@@
A particularly intriguing feature of massive perturbations is the possibility of arbitrarily long-lived modes when the field mass approaches certain critical values Ohashi and Sakagami (2004); Konoplya and Zhidenko (2005). This phenomenon has been observed for different spins of the perturbing field Konoplya (2006); Fernandes et al. (2022); Konoplya and Zhidenko (2018); Percival and Dolan (2020), various black-hole backgrounds Zhidenko (2006); Zinhailo (2018); Churilova (2020); Bolokhov (2024d). The presence of the mass term also qualitatively modifies the late-time decay of perturbations: instead of the usual power-law tails characteristic of massless fields, the signal exhibits oscillatory late-time behavior Jing (2005); Koyama and Tomimatsu (2002); Moderski and Rogatko (2001); Rogatko and Szyplowska (2007); Koyama and Tomimatsu (2001b, a); Gibbons et al. (2008); Gibbons and Rogatko (2008). In certain situations, an effective mass term may also appear dynamically, for example, when a massless field propagates in the vicinity of a black hole immersed in a magnetic field Konoplya and Fontana (2008); Konoplya (2008); Wu and Xu (2015). At the same time, the existence of arbitrarily long-lived modes is not universal, and there are examples where such modes do not occur Zinhailo (2024); Konoplya (2006). However, the behavior of massive perturbations and their associated observational characteristics have not yet been systematically studied for regular black holes arising in four-dimensional quasi-topological gravity.
@@ II Regular black-hole geometries in quasi-topological gravity
Non-polynomial quasi-topological gravity (NP-QTG) represents a class of modified gravitational theories in which higher-curvature invariants are combined in such a way that the resulting dynamics remains well behaved in symmetric backgrounds Bueno et al. (2025b, a, c); Frolov et al. (2025). The action of the theory can be written schematically as Bueno et al. (2026); Borissova and Carballo-Rubio (2026)
@@ IV.1 WKB approximation
In this work, we make use of high-order WKB approximations following the formulations of Refs. Matyjasek and Telecka (2019); Matyjasek and Opala (2017). For backgrounds that admit analytic treatment, we employ the 14th- and 16th-order expressions. A useful practical criterion of reliability is the stability of the result with respect to the WKB order: when several neighboring orders produce nearly the same quasinormal frequency, the expected uncertainty is typically comparable to the difference between these consecutive approximations. At the same time, it should be kept in mind that the WKB expansion represents an asymptotic series, so that increasing the order does not necessarily lead to a strictly monotonic improvement of accuracy. The WKB technique has been widely elaborated and employed in the analysis of black hole perturbations (see, for instance, Konoplya and Zhidenko (2007); Malik (2024); Hamil and Lütfüoğlu (2025); Skvortsova (2024a); Han and Gwak (2026); Konoplya and Stashko (2025); Malik (2025c); Kodama et al. (2010); Lütfüoğlu et al. (2026b); Arbelaez (2026); Konoplya and Zhidenko (2023); Bolokhov (2024b); Lütfüoğlu et al. (2026a); Konoplya and Zhidenko (2010); Kanti et al. (2006); Bolokhov (2024a); Lütfüoğlu (2025a); Skvortsova (2024b); Bolokhov (2025b); Lütfüoğlu (2026); Stuchlík and Zhidenko (2025b); Bolokhov and Skvortsova (2025c)), and therefore we refrain from repeating a detailed description of the method here.
@@ IV.2 Time-domain integration
The resulting time-domain profiles typically consist of three stages: an initial transient determined by the chosen initial data, a period of exponentially damped oscillations dominated by QNMs, and finally the late-time tail. The quasinormal frequencies can be extracted from the intermediate stage using fitting procedures such as the Prony method, which decomposes the signal into a superposition of damped exponentials.
The time-domain integration approach has been widely applied in the study of black hole perturbations. It provides a robust way to extract the dominant quasinormal mode and to reveal possible dynamical instabilities from the temporal evolution of perturbations. Owing to these advantages, the method has been implemented in numerous investigations (see, for example, Konoplya and Zhidenko (2014b); Skvortsova (2024c); Dubinsky (2026b); Ishihara et al. (2008); Konoplya and Molina (2005); Stuchlík and Zhidenko (2025a); Dubinsky (2024); Abdalla et al. (2005); Lütfüoğlu et al. (2025b); Cuyubamba et al. (2016); Konoplya and Zhidenko (2014a); Lütfüoğlu et al. (2025a); Konoplya et al. (2023); Dubinsky (2025a); Skvortsova (2025a); Lütfüoğlu (2025b); Dubinsky (2025b); Bolokhov and Skvortsova (2025b)), demonstrating its effectiveness for determining the fundamental oscillation frequency and diagnosing instabilities.
@@
Table 3: Event horizon position ($r_{0}$), Hawking temperature ($T_{H}$), shadow radius ($R_{s}$), Lyapunov exponent ($\lambda$), ISCO frequency ($\Omega_{ISCO}$) and binding energy (BE) for the regular black hole I ($M=1$).
@@
Table 4: Event horizon position ($r_{0}$), Hawking temperature ($T_{H}$), shadow radius ($R_{s}$), Lyapunov exponent ($\lambda$), ISCO frequency ($\Omega_{ISCO}$) and binding energy (BE) for the regular black hole ($M=1$).
@@ VI.1 Photon sphere and shadow
$\lambda^{2}=-\frac{1}{P}\frac{d^{2}P}{dr_{*}^{2}}\Bigg|_{r=r_{m}}.$ (31)
@@ VII Conclusions
Overall, the considered regular black-hole geometries preserve the main qualitative properties of black-hole perturbations and geodesic motion while introducing moderate quantitative deviations. The results demonstrate that massive scalar perturbations in these spacetimes naturally approach the long-lived quasi-resonant regime, providing a characteristic signature of massive-field dynamics around regular black holes.

### 16 Apr 2026 v2

@@ Abstract
We investigate massive scalar perturbations and several characteristics of particle motion in the spacetime of regular black holes arising in four-dimensional quasi-topological gravity. Quasinormal modes are computed using high-order WKB approximations with Padé resummation and verified through time-domain integration. For moderate values of the scalar-field mass, the time-domain profiles confirm the WKB results with excellent accuracy. As the mass increases, the damping rate decreases substantially, indicating the approach to the quasi-resonant regime of long-lived modes. For sufficiently large masses, the late-time signal becomes dominated by oscillatory power-law tails, which mask the quasi-resonant mode in the time-domain profile. In addition, we analyze photon motion and circular geodesics, including the photon-sphere radius, shadow size, Lyapunov exponent, and ISCO characteristics. These quantities exhibit only moderate deviations from their Schwarzschild values, unlike the Hawking temperature of the black hole.
@@ I Introduction
One of the recent findings in this area is the construction of regular black holes in four-dimensional non-polynomial quasi-topological gravity Borissova and Carballo-Rubio (2026); Bueno et al. (2026a). Within this framework, black-hole solutions can be obtained analytically. The resulting geometries are asymptotically flat and free of curvature singularities, reproducing well-known regular metrics without introducing additional matter sources. A regularization parameter usually controls the deviation from the Schwarzschild spacetime and is typically localized in the near-horizon region.
Understanding how such geometrically regular black holes manifest themselves through observable phenomena requires analyzing both wave dynamics and particle motion in these spacetimes. One of the most important probes of black-hole geometry is provided by quasinormal modes (QNMs), which describe the characteristic oscillations of perturbed black-holes Konoplya and Zhidenko (2011); Bolokhov and Skvortsova (2025c); Kokkotas and Schmidt (1999). These modes dominate the ringdown stage of gravitational-wave signals produced in black-hole mergers and therefore play a key role in extracting information about the underlying black-hole spacetime geometry Abbott and others (2016, 2017, 2020). At the same time, properties of geodesic motion, such as the angular velocity at the innermost stable circular orbit (ISCO), the Lyapunov exponent governing the instability of photon orbits, the binding energy of particles, and the radius of the black-hole shadow, provide complementary observational diagnostics that directly probe the structure of the gravitational field in the strong-curvature regime.
@@
A particularly intriguing feature of massive perturbations is the possibility of arbitrarily long-lived modes when the field mass approaches certain critical values Ohashi and Sakagami (2004); Konoplya and Zhidenko (2005). This phenomenon has been observed for different spins of the perturbing field Konoplya (2006); Fernandes et al. (2022); Konoplya and Zhidenko (2018); Percival and Dolan (2020), various black-hole backgrounds Zhidenko (2006); Zinhailo (2018); Churilova (2020); Bolokhov (2024d). The presence of the mass term also qualitatively modifies the late-time decay of perturbations: instead of the usual power-law tails characteristic of massless fields, the signal exhibits oscillatory late-time behavior Jing (2005); Koyama and Tomimatsu (2002); Moderski and Rogatko (2001); Rogatko and Szyplowska (2007); Koyama and Tomimatsu (2001b, a); Gibbons et al. (2008); Gibbons and Rogatko (2008). In certain situations, an effective mass term may also appear dynamically, for example, when a massless field propagates in the vicinity of a black hole immersed in a magnetic field Konoplya and Fontana (2008); Konoplya (2008); Wu and Xu (2015). At the same time, the existence of arbitrarily long-lived modes is not universal, and there are examples where such modes do not occur Zinhailo (2024); Konoplya (2006). However, the behavior of massive perturbations and their associated observational characteristics have not yet been studied for regular black holes arising in four-dimensional quasi-topological gravity.
@@ II Regular black-hole geometries in quasi-topological gravity
Non-polynomial quasi-topological gravity (NP-QTG) represents a class of modified gravitational theories in which higher-curvature invariants are combined in such a way that the resulting dynamics remains well behaved in symmetric backgrounds Bueno et al. (2025b, a, 2026b); Frolov et al. (2025). The action of the theory can be written schematically as Bueno et al. (2026a); Borissova and Carballo-Rubio (2026)
@@
While there exists a large freedom in the choice of generating functions leading to a wide variety of black hole solutions—many of which can be regular—a few particularly representative metrics, such as those of Hayward Hayward (2006), Dymnikova Dymnikova (1992); Dymnikova and Khlopov (2015), Bardeen Bardeen (1968) and related models, have been extensively investigated in the literature Fernando and Correa (2012); Saleh et al. (2018); Konoplya (2023); Konoplya et al. (2023b); Lütfüoğlu et al. (2026b, c); Konoplya et al. (2022); Lütfüoğlu (2026a). In the present work, we consider a new regular black hole metric possessing a set of desirable properties, which will be specified below.
We demonstrate that, when analyzed alongside previously studied cases, this solution exhibits a number of common features at the level of observable quantities. In particular, higher-curvature corrections tend to produce qualitatively similar effects across different models. This universality enhances the significance of our results: although specific black hole constructions may differ, one can expect broadly similar spectral and orbital characteristics for black holes modified by higher-curvature contributions.
@@ IV.1 WKB approximation
In this work, we make use of high-order WKB approximations following the formulations of Refs. Matyjasek and Telecka (2019); Matyjasek and Opala (2017). For backgrounds that admit analytic treatment, we employ the 14th- and 16th-order expressions. A useful practical criterion of reliability is the stability of the result with respect to the WKB order: when several neighboring orders produce nearly the same quasinormal frequency, the expected uncertainty is typically comparable to the difference between these consecutive approximations. At the same time, it should be kept in mind that the WKB expansion represents an asymptotic series, so that increasing the order does not necessarily lead to a strictly monotonic improvement of accuracy. The WKB technique has been widely elaborated and employed in the analysis of black hole perturbations (see, for instance, Konoplya and Zhidenko (2007); Malik (2024); Hamil and Lütfüoğlu (2025); Skvortsova (2024a); Han and Gwak (2026); Konoplya and Stashko (2025); Malik (2025c); Kodama et al. (2010); Lütfüoğlu et al. (2026c); Arbelaez (2026); Konoplya and Zhidenko (2023); Bolokhov (2024b); Lütfüoğlu et al. (2026a); Konoplya and Zhidenko (2010); Kanti et al. (2006); Bolokhov (2024a); Lütfüoğlu (2025a); Skvortsova (2024b); Bolokhov (2025b); Lütfüoğlu (2026b); Stuchlík and Zhidenko (2025b); Bolokhov and Skvortsova (2026)), and therefore we refrain from repeating a detailed description of the method here.
@@ IV.2 Time-domain integration
In our implementation of the time-domain integration, we follow the characteristic scheme of Gundlach, Price, and Pullin Gundlach et al. (1994), discretizing the wave equation on a uniform grid in the light-cone coordinates $(u,v)$ according to Eq. (18). The evolution is initiated by specifying the field on the initial null surfaces, where we choose a Gaussian wave packet of the form $\Psi(u=u_{0},v)=\exp[-(v-v_{c})^{2}/(2\sigma^{2})]$ and $\Psi(u,v=v_{0})=0$, with $v_{c}$ determining the initial location of the pulse and $\sigma$ its width. The overall normalization of the initial data is arbitrary due to the linearity of the perturbation equation; throughout the evolution we therefore work with dimensionless amplitudes, so that the vertical scale in Figs. 3 and 4 reflects the relative decay of the signal rather than an absolute physical normalization. The resulting time-domain profile consists of an initial transient stage, followed by a period of exponentially damped oscillations dominated by QNMs, and finally a late-time tail, in agreement with the general expectations for black-hole perturbations. The quasinormal frequencies are extracted from the intermediate stage using the Prony method, which approximates the signal as a superposition of damped exponentials $\Psi(t)=\sum_{k}C_{k}e^{-i\omega_{k}t}$ and determines the complex frequencies $\omega_{k}$ via a linear fitting procedure (see, London et al. (2014); Bolokhov (2026); Lütfüoğlu et al. (2025b); Konoplya and Zhidenko (2022) for recent examples). In practice, we select a time window where the ringing stage is clearly established and the contamination from both the initial burst and the late-time tail is negligible; the stability of the extracted frequencies with respect to variations of this window provides an estimate of the numerical uncertainty.
The time-domain integration approach has been widely applied in the study of black hole perturbations. It provides a robust way to extract the dominant quasinormal mode and to reveal possible dynamical instabilities from the temporal evolution of perturbations. Owing to these advantages, the method has been implemented in numerous investigations (see, for example, Konoplya (2021); Konoplya and Zhidenko (2014b); Skvortsova (2024c); Dubinsky (2026b); Ishihara et al. (2008); Konoplya and Molina (2005); Stuchlík and Zhidenko (2025a); Dubinsky (2024); Abdalla et al. (2005); Cuyubamba et al. (2016); Konoplya and Zhidenko (2014a); Lütfüoğlu et al. (2025a); Konoplya et al. (2023a); Dubinsky (2025a); Skvortsova (2025a); Lütfüoğlu (2025b); Dubinsky (2025b); Bolokhov and Skvortsova (2025b)), demonstrating its effectiveness for determining the fundamental oscillation frequency and diagnosing instabilities.
@@
Table 3: Event horizon position ($r_{0}$), Hawking temperature ($T_{H}$), photon-sphere radius ($r_{m}$), shadow radius ($R_{s}$), Lyapunov exponent ($\lambda$), ISCO frequency ($\Omega_{ISCO}$) and binding energy (BE) for the regular black hole I ($M=1$).
@@
Table 4: Event horizon position ($r_{0}$), Hawking temperature ($T_{H}$), photon-sphere radius ($r_{m}$), shadow radius ($R_{s}$), Lyapunov exponent ($\lambda$), ISCO frequency ($\Omega_{ISCO}$) and binding energy (BE) for the regular black hole ($M=1$).
@@ V Quasinormal modes
Although gravitational perturbations generally differ from those of test fields, numerous studies have shown that their spectra share qualitatively similar features, especially at moderate and high multipole numbers. This is because the dominant contribution to the effective potential is given by the centrifugal term $\ell(\ell+1)/r^{2}$, which quickly outweighs subleading corrections, so that the test-field approximation frequently captures the main characteristics of the spectrum.
@@ VI.1 Photon sphere and shadow
$\lambda^{2}=-\frac{1}{2P}\frac{d^{2}P}{dr_{*}^{2}}\Bigg|_{r=r_{m}}.$ (31)
@@ VII Conclusions
Overall, we observe that observables related to particle motion and the least damped QNMs exhibit only a mild dependence on the coupling parameter responsible for regularization. In contrast, quantities determined by the near-horizon geometry, such as the Hawking temperature, can vary significantly with the coupling. This behavior is a common feature of various regular black hole models, including those of Hayward, Dymnikova, and Bardeen. It is apparently related to the fact that the least damped modes, as well as key characteristics of particle motion, are predominantly governed by the geometry in the region between the photon sphere and the innermost stable circular orbit, i.e., at a finite distance from the event horizon. At the same time, these regular black hole solutions exhibit substantial deviations from the Schwarzschild geometry in the immediate vicinity of the horizon, while rapidly approaching the Schwarzschild limit at larger radial distances. The results demonstrate that massive scalar perturbations in these spacetimes naturally approach the long-lived quasi-resonant regime, providing a characteristic signature of massive-field dynamics around regular black holes.